The student council is hosting a drawing to raise money for scholarships. They are selling tickets for $6 each and will sell 900 tickets. There is one $3,000 grand prize, two $300 second prizes, and fifteen $10 third prizes. You just bought a ticket. Find the expected value for your profit. Round to the nearest cent.


$_______

Answer by math_tutor2020(3835)


Method 1

X = net gain

Define the four events:
A = win the 1st prize $3000
B = win the 2nd prize $300
C = win the 3rd prize $10
D = no winnings

If you paid $6 and win $3000, then the net winnings is 3000-6 = 2994. The chances of this happening is 1 out of 900, aka 1/900.

If you paid $6 and win $300, then the net winnings is 300-6 = 294. The chances of this happening is 2/900.

If you paid $6 and win $10, then the net winnings is 10-6 = 4. The chances of this happening is 15/900.

There are 1+2+15 = 18 winning tickets.
That means there are 900-18 = 882 non-winning tickets.
The chances of not winning any money is 882/900.

Probability Distribution Table:
EventXP(X)
A29941/900
B2942/900
C415/900
D-6882/900

where X = net gain


Form a fourth column X*P(X) to multiply each X value with its corresponding probability P(X)
EventXP(X)X*P(X)
A29941/9002994/900
B2942/900588/900
C415/90060/900
D-6882/900-5292/900

Add up the results in the X*P(X) column to get the expected value.
2994/900 + 588/900 + 60/900 + (-5292/900)
(2994+588+60-5292)/900
-1650/900
-1.83333333333333
-1.83

You should expect to lose approximately $1.83 per ticket on average.

Answer: -1.83
Don't forget about the negative sign to indicate a loss.


========================================================

Method 2

Let's say that you purchased all 900 tickets.

1 ticket costs 6 dollars.
900 tickets cost 900*6 = 5400 dollars.
So far you are down $5400.

But the one grand prize $3000, two $300 prizes, and fifteen $10 prizes add up to:
1*3000+2*300+15*10 = 3750 dollars

The net change is
amountIn - amountOut = 3750-5400 = -1650
Meaning you are down $1650

The loss per ticket is: 1650/900 = 1.83 after rounding to two decimal places.

Answer: -1.83
Don't forget about the negative sign to indicate a loss.

Method 2 is much easier to deal with, since we avoid messy fractions.
However, method 1 is a standard approach that many math textbooks and teachers use.
It's good to know how to do both methods.